OpenStudy (emmaleelooney):

hheeeellpppp! The legs of an isosceles triangle have lengths x + 1 and -x + 7 . The base has length 3x - 3. What is the length of the base? A. 4 B. 6 C. 3 D. cannot be determined

Directrix (directrix):

The legs of an isosceles triangle are congruent. First step, Solve this equation for x: x + 1 = -x + 7

Directrix (directrix):

@emmaleelooney

syed98 (syedmohammed98):

i think it is 6 as u solve the equation x+1 = -x +7 u get x=3 so 3x-3-----> 3(3)-3----> gives you 6 :)